OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ x ^{2} - 16 }{ x ^{2} - 4x }\]

OpenStudy (anonymous):

=> \[\frac{ (x+4)(x-4) }{ x(x-4)? }\]

OpenStudy (anonymous):

=> \[\frac{ x+4 }{ x } => 1 + \frac{ 4 }{ x }\]